# COMET: Co-Optimization of a CNN Model using Efficient-Hardware OBC Techniques

### v1

@@ V-B Comparison with State-of-the-Art CNNs
  • [15] reports 19.20 GOP/s using 72 DSPs, 80 BRAMs, 17,420 LUTs, and 10,195 FFs at 100 MHz. With an ENS of $\approx$21,024 and an AEP of 0.913 ops/cycle/kENS, this design exhibits better normalized efficiency than [16], but DSPs and BRAMs remain underutilized for low-precision operations.
@@
  • It is worth noting that [15] does not provide power consumption results, limiting a complete efficiency assessment. While its reported AEP appears higher than other works, the design occupies significantly more FFs than all the proposed variants, making it less favorable in terms of overall resource efficiency.
@@ VI Conclusion
In this work, efficient CNN implementations based on the proposed hardware OBC techniques have been presented. By formulating CNN inference in OBC form for both inputs and weights, bit-width asymmetry has been effectively exploited and its influence on hardware efficiency analyzed. Furthermore, OBC-based LUT implementations—parallel, shared, split, and hybrid—have been developed and evaluated to highlight their trade-offs in performance, resource usage, and energy efficiency. These optimizations have been integrated into an OBC-GEMM core using the im2col transformation, enabling streamlined and resource-aware CNN execution. These methods have been encapsulated within the COMET framework, providing a unified platform for co-optimized CNN acceleration. Experimental results on a fixed-point modified LeNet-5 have demonstrated substantial reductions in resource utilization compared to state-of-the-art CNNs based on LeNet-5 [16]. By leveraging FPGA LUT fabric and eliminating reliance on DSPs and BRAMs in the compute datapath, the proposed approach offers a scalable solution for lightweight CNN deployment on edge devices. Overall, this work establishes OBC-based computation as a promising and extensible paradigm for co-optimized CNN acceleration across diverse edge AI applications.

### v2

@@ V-B Comparison with State-of-the-Art CNNs
  • [15] reports 19.20 GOP/s using 72 DSPs, 80 BRAMs, 17,420 LUTs, and 10,195 FFs at 100 MHz. With an ENS of $\approx$21,024 and an AEP of 0.913 ops/cycle/kENS, this design exhibits better normalized efficiency than [16], but involve DSPs and BRAMs remain underutilized for low-precision operations.
@@
  • Although [15] reports a higher AEP than our proposed variants, its evaluation does not include power consumption results, preventing a complete assessment of overall efficiency. Moreover, the design requires substantially more flip-flops (FFs), which directly contribute to higher dynamic power and area usage, thereby reducing its practical resource efficiency compared with our implementations.
@@ VI Conclusion
We presented an efficient CNN implementations based on the proposed hardware OBC techniques. By formulating CNN inference in OBC form for both inputs and weights, bit-width asymmetry has been effectively exploited and its influence on hardware efficiency analyzed. Furthermore, OBC-based LUT implementations—parallel, shared, split, and hybrid—have been developed and evaluated to highlight their trade-offs in performance, resource usage, and energy efficiency. These optimizations have been integrated into an OBC-GEMM core using the im2col transformation, enabling streamlined and resource-aware CNN execution. These methods have been encapsulated within the COMET framework, providing a unified platform for co-optimized CNN acceleration. Experimental results on a fixed-point modified LeNet-5 have demonstrated substantial reductions in resource utilization compared to state-of-the-art CNNs based on LeNet-5 [16]. By leveraging FPGA LUT fabric and eliminating reliance on DSPs and BRAMs in the compute datapath, the proposed approach offers a scalable solution for lightweight CNN deployment on edge devices. Overall, this work establishes OBC-based computation as a promising and extensible paradigm for co-optimized CNN acceleration across diverse edge AI applications.